# Know “No” Better: A Data-Driven Approach for Enhancing Negation Awareness in CLIP

### v1

@@ E.3 Text-to-Image Generation
Prompt Question Prompt Question
a man not wearing a hat Is the man wearing a hat? a woman not wearing a mask Is the woman wearing a mask?
a dog not running Is the dog running? a bird not flying Is the bird flying?
a not white cat Is the cat white? a not blue sphere Is the sphere blue?
@@
a room with no window Is there a window in the room? a dog without a collar Does the dog have a collar?
a cat with no whiskers Does the cat have whiskers? a child not holding a toy Is the child holding a toy?
@@
a park with no benches Are there benches in the park? a bird not perched on a branch Is the bird perched on a branch?
@@
a woman without glasses Is the woman wearing glasses? a table with no chairs around Are there chairs around the table?
a car not parked in the driveway Is the car parked in the driveway? a beach without any umbrellas Are there umbrellas on the beach?
a man with no hat Is the man wearing a hat? a road not crowded with cars Is the road crowded with cars?
@@
a garden with no flowers Are there flowers in the garden? a person not holding an umbrella Is the person holding an umbrella?
@@
a lake with no boats Are there boats on the lake? a house without a roof Does the house have a roof?
a tree not in bloom Is the tree in bloom? a mountain with no snow Is there snow on the mountain?
a room without furniture Is there furniture in the room? a dog not barking Is the dog barking?
@@
a street with no people Are there people on the street? a kitchen without any food Is there food in the kitchen?
a cup not filled with coffee Is the cup filled with coffee? a forest with no animals Are there animals in the forest?
a phone not on the table Is the phone on the table? a desk without a computer Is there a computer on the desk?
a man not wearing shoes Is the man wearing shoes? a restaurant with no tables Are there tables in the restaurant?
@@
a city skyline with no skyscrapers Are there skyscrapers in the city skyline? a field without crops Are there crops in the field?
@@
a woman not smiling Is the woman smiling? a living room with no couch Is there a couch in the living room?
@@
a car without wheels Does the car have wheels? a stadium with no spectators Are there spectators in the stadium?
@@
a road with no signs Are there signs on the road? a child not wearing shoes Is the child wearing shoes?
@@
a bridge without railings Does the bridge have railings? a river with no fish Are there fish in the river?
@@
a sky without clouds Are there clouds in the sky? a cup with no handle Does the cup have a handle?
a playground with no swings Are there swings in the playground? a man not wearing a tie Is the man wearing a tie?
@@
a building without windows Does the building have windows? a book with no cover Does the book have a cover?
@@
a shop with no customers Are there customers in the shop? a garden without any trees Are there trees in the garden?
@@
a bike not leaning against a wall Is the bike leaning against a wall? a stage with no performers Are there performers on the stage?
@@
a train station with no trains Are there trains at the train station? a museum without exhibits Are there exhibits in the museum?
@@
a shelf with no books Are there books on the shelf? a restaurant not serving food Is the restaurant serving food?
@@
a person with no backpack Does the person have a backpack? a market without any vendors Are there vendors in the market?
@@
a room not filled with light Is the room filled with light? a path with no signs Are there signs on the path?
@@
a school without students Are there students in the school? a car with no headlights Does the car have headlights?
a cat without a tail Does the cat have a tail? a person not holding a bag Is the person holding a bag?
@@
a forest with no leaves Are there leaves in the forest? a house with no doors Does the house have doors?
a chair not facing the table Is the chair facing the table? a bird not singing Is the bird singing?
a beach without sand Is there sand on the beach? a dog not playing fetch Is the dog playing fetch?
a wall with no decorations Are there decorations on the wall? a sidewalk with no pedestrians Are there pedestrians on the sidewalk?
a man not reading a book Is the man reading a book? a classroom with no desks Are there desks in the classroom?
@@
a street with no lights Are there lights on the street? a yard without grass Is there grass in the yard?
a riverbank with no trees Are there trees on the riverbank? a cat not purring Is the cat purring?
@@
a boat with no sails Does the boat have sails? a woman not holding a purse Is the woman holding a purse?
a stadium with no players Are there players in the stadium? a sky with no stars Are there stars in the sky?
a store without shelves Are there shelves in the store? a man not holding a briefcase Is the man holding a briefcase?
@@
a city without buildings Are there buildings in the city? a painting without colors Does the painting have colors?
@@
a road without any turns Are there turns on the road? a lawn with no flowers Are there flowers on the lawn?
a dog not fetching a ball Is the dog fetching a ball? a bridge without any lights Are there lights on the bridge?
a car with no passengers Are there passengers in the car? a garden with no vegetables Are there vegetables in the garden?
@@
a child not drinking milk Is the child drinking milk? a person without a shadow Does the person have a shadow?
@@
a tree with no leaves Does the tree have leaves? a bus stop without a bench Is there a bench at the bus stop?
@@
a train without passengers Are there passengers on the train? a cafe with no tables Are there tables in the cafe?
@@
a photo with no people Are there people in the photo? a street without any trees Are there trees on the street?
a river not flowing Is the river flowing? a mountain with no trails Are there trails on the mountain?
@@
a path without any footprints Are there footprints on the path? a field with no animals Are there animals in the field?
a building with no entrance Is there an entrance to the building?
@@
For generating negation-inclusive prompts, we utilized ChatGPT [29] to construct 107 prompts. To evaluate whether the generated images accurately reflected the negation-related aspects of the prompts, we also employed ChatGPT to create corresponding questions. These questions were designed such that the expected answer from the MLLM evaluator would be “no”. The complete list of prompts and their corresponding questions is provided in  Tab. 11.
In SD-1.4 [34], we replaced its CLIP ViT-L/14 text encoder with our fine-tuned negation-aware text encoder of the same architecture. For SDXL, which employs two text encoders (CLIP ViT-L/14 and CLIP ViT-BigG/14), we substituted both encoders with our fine-tuned versions.
@@ E.4 Referring Image Segmentation
For our experiments in referring image segmentation in Sec. 5.2, we utilized the publicly available weight of CLIPSeg [26] trained on the PhraseCut [44] dataset as our baseline. Without any additional training, we replaced the text encoder in the CLIPSeg architecture with our fine-tuned CLIP ViT-B/16 text encoder.
@@
Appendix F Additional Qualitative Results

### v2

@@ E.3 T2I Generation
Prompt Question 1 Question 2 Prompt Question 1 Question 2
a man not wearing a hat Is this a man? Is the man wearing a hat? a woman not wearing a mask Is this a woman? Is the woman wearing a mask?
a dog not running Is this a dog? Is the dog running? a bird not flying Is this a bird? Is the bird flying?
a not white cat Is this a cat? Is the cat white? a not blue sphere Is this a sphere? Is the sphere blue?
@@
a room with no window Is this a room? Is there a window in the room? a dog without a collar Is this a dog? Does the dog have a collar?
a cat with no whiskers Is this a cat? Does the cat have whiskers? a child not holding a toy Is this a child? Is the child holding a toy?
@@
a park with no benches Is this a park? Are there benches in the park? a bird not perched on a branch Is this a bird? Is the bird perched on a branch?
@@
a woman without glasses Is this a woman? Is the woman wearing glasses? a table with no chairs around Is this a table? Are there chairs around the table?
a car not parked in the driveway Is this a car? Is the car parked in the driveway? a beach without any umbrellas Is this a beach? Are there umbrellas on the beach?
a man with no hat Is this a man? Is the man wearing a hat? a road not crowded with cars Is this a road? Is the road crowded with cars?
@@
a garden with no flowers Is this a garden? Are there flowers in the garden? a person not holding an umbrella Is this a person? Is the person holding an umbrella?
@@
a lake with no boats Is this a lake? Are there boats on the lake? a house without a roof Is this a house? Does the house have a roof?
a tree not in bloom Is this a tree? Is the tree in bloom? a mountain with no snow Is this a mountain? Is there snow on the mountain?
a room without furniture Is this a room? Is there furniture in the room? a dog not barking Is this a dog? Is the dog barking?
@@
a street with no people Is this a street? Are there people on the street? a kitchen without any food Is this a kitchen? Is there food in the kitchen?
a cup not filled with coffee Is this a cup? Is the cup filled with coffee? a forest with no animals Is this a forest? Are there animals in the forest?
a phone not on the table Is this a phone? Is the phone on the table? a desk without a computer Is this a desk? Is there a computer on the desk?
a man not wearing shoes Is this a man? Is the man wearing shoes? a restaurant with no tables Is this a restaurant? Are there tables in the restaurant?
@@
a city skyline with no skyscrapers Is this a city skyline? Are there skyscrapers in the city skyline? a field without crops Is this a field? Are there crops in the field?
@@
a woman not smiling Is this a woman? Is the woman smiling? a living room with no couch Is this a living room? Is there a couch in the living room?
@@
a car without wheels Is this a car? Does the car have wheels? a stadium with no spectators Is this a stadium? Are there spectators in the stadium?
@@
a road with no signs Is this a road? Are there signs on the road? a child not wearing shoes Is this a child? Is the child wearing shoes?
@@
a bridge without railings Is this a bridge? Does the bridge have railings? a river with no fish Is this a river? Are there fish in the river?
@@
a sky without clouds Is this a sky? Are there clouds in the sky? a cup with no handle Is this a cup? Does the cup have a handle?
a playground with no swings Is this a playground? Are there swings in the playground? a man not wearing a tie Is this a man? Is the man wearing a tie?
@@
a building without windows Is this a building? Does the building have windows? a book with no cover Is this a book? Does the book have a cover?
@@
a shop with no customers Is this a shop? Are there customers in the shop? a garden without any trees Is this a garden? Are there trees in the garden?
@@
a bike not leaning against a wall Is this a bike? Is the bike leaning against a wall? a stage with no performers Is this a stage? Are there performers on the stage?
@@
a train station with no trains Is this a train station? Are there trains at the train station? a museum without exhibits Is this a museum? Are there exhibits in the museum?
@@
a shelf with no books Is this a shelf? Are there books on the shelf? a restaurant not serving food Is this a restaurant? Is the restaurant serving food?
@@
a person with no backpack Is this a person? Does the person have a backpack? a market without any vendors Is this a market? Are there vendors in the market?
@@
a room not filled with light Is this a room? Is the room filled with light? a path with no signs Is this a path? Are there signs on the path?
@@
a school without students Is this a school? Are there students in the school? a car with no headlights Is this a car? Does the car have headlights?
a cat without a tail Is this a cat? Does the cat have a tail? a person not holding a bag Is this a person? Is the person holding a bag?
@@
a forest with no leaves Is this a forest? Are there leaves in the forest? a house with no doors Is this a house? Does the house have doors?
a chair not facing the table Is this a chair? Is the chair facing the table? a bird not singing Is this a bird? Is the bird singing?
a beach without sand Is this a beach? Is there sand on the beach? a dog not playing fetch Is this a dog? Is the dog playing fetch?
a wall with no decorations Is this a wall? Are there decorations on the wall? a sidewalk with no pedestrians Is this a sidewalk? Are there pedestrians on the sidewalk?
a man not reading a book Is this a man? Is the man reading a book? a classroom with no desks Is this a classroom? Are there desks in the classroom?
@@
a street with no lights Is this a street? Are there lights on the street? a yard without grass Is this a yard? Is there grass in the yard?
a riverbank with no trees Is this a riverbank? Are there trees on the riverbank? a cat not purring Is this a cat? Is the cat purring?
@@
a boat with no sails Is this a boat? Does the boat have sails? a woman not holding a purse Is this a woman? Is the woman holding a purse?
a stadium with no players Is this a stadium? Are there players in the stadium? a sky with no stars Is this a sky? Are there stars in the sky?
a store without shelves Is this a store? Are there shelves in the store? a man not holding a briefcase Is this a man? Is the man holding a briefcase?
@@
a city without buildings Is this a city? Are there buildings in the city? a painting without colors Is this a painting? Does the painting have colors?
@@
a road without any turns Is this a road? Are there turns on the road? a lawn with no flowers Is this a lawn? Are there flowers on the lawn?
a dog not fetching a ball Is this a dog? Is the dog fetching a ball? a bridge without any lights Is this a bridge? Are there lights on the bridge?
a car with no passengers Is this a car? Are there passengers in the car? a garden with no vegetables Is this a garden? Are there vegetables in the garden?
@@
a child not drinking milk Is this a child? Is the child drinking milk? a person without a shadow Is this a person? Does the person have a shadow?
@@
a tree with no leaves Is this a tree? Does the tree have leaves? a bus stop without a bench Is this a bus stop? Is there a bench at the bus stop?
@@
a train without passengers Is this a train? Are there passengers on the train? a cafe with no tables Is this a cafe? Are there tables in the cafe?
@@
a photo with no people Is this a photo? Are there people in the photo? a street without any trees Is this a street? Are there trees on the street?
a river not flowing Is this a river? Is the river flowing? a mountain with no trails Is this a mountain? Are there trails on the mountain?
@@
a path without any footprints Is this a path? Are there footprints on the path? a field with no animals Is this a field? Are there animals in the field?
a building with no entrance Is this a building? Is there an entrance to the building?
@@
For generating negation-inclusive prompts, we utilized ChatGPT [32] to construct 107 prompts. To evaluate whether the generated images accurately reflected the given prompts, we employed ChatGPT to create two corresponding questions for each prompt. The first question was designed to assess whether the subject of the prompt was correctly generated, with the expected answer being “yes.” The second question aimed to verify whether the negation-related aspect was properly removed, with the expected answer being “no.” For example, for the prompt “a man not wearing a hat,” the two generated questions were “Is this a man?” and “Is the man wearing a hat?”. The complete list of prompts and their corresponding questions is provided in  Tab. 11.
In SD-1.4 [37], we replaced its CLIP ViT-L/14 text encoder with CoN-CLIP text encoder and our fine-tuned negation-aware text encoder of the same architecture. For SDXL, which employs two text encoders (CLIP ViT-L/14 and CLIP ViT-BigG/14), we substituted both encoders with our fine-tuned versions.
@@ E.4 Referring Image Segmentation
For our experiments in referring image segmentation in Sec. 5.2, we utilized the publicly available weight of CLIPSeg [28] trained on the PhraseCut [48] dataset as our baseline. Without any additional training, we replaced the text encoder in the CLIPSeg architecture with CoN-CLIP ViT-B/16 text encoder and our fine-tuned CLIP ViT-B/16 text encoder.
@@
Appendix F Expanded Discussion on T2I Generation
F.1 Methods
In T2I generation, research has emerged focusing on the removal of unwanted concepts from generated images. [8, 22, 49]. These methods typically preprocess text prompts by identifying and removing negation-related components, reformulating the prompt to exclude negation. For example, given the text prompt “a panda in a forest without flowers,” LMD [22] uses an LLM is used to identify “flowers” as the negated object and remove it, resulting in a final layout-based prompt that retains only “a panda” and “a forest.”
In contrast, our approach directly uses the original prompt “a panda in a forest without flowers” without any preprocessing, allowing the T2I model to process negation as part of the input text. This provides a different perspective on handling negation, where the model learns to interpret negation within natural language rather than relying on preprocessing to modify the prompt.
F.2 Evaluation
Existing methods evaluate negation comprehension in T2I models using the LMD [22] Negation benchmark, which consists of 10 prompts structured as “a realistic photo of a scene without [object]” and uses object detectors to verify object absence.
Table 12: Results on LMD Negation benchmark.
Method LMD Negation
SD 20%
SD w/ NegationCLIP text encoder 94%
SD + LMD [22] 100%
We provide evaluation results of NegationCLIP on the LMD Negation benchmark in Tab. 12. Standard SD struggles significantly with negation, achieving only 20% accuracy, indicating that it often fails to remove the negated object. In contrast, simply replacing SD’s text encoder with NegationCLIP’s text encoder improves accuracy to 94%, successfully removing the negated object in nearly all cases. Unsurprisingly, applying LMD [22] method to SD yields perfect performance, as it heuristically removes the negation-related part from the final prompt, ensuring that the object does not appear in the generated image.
However, while the LMD Negation benchmark effectively evaluates object removal, it does not account for negation beyond object presence or absence, such as actions and attributes. While it can determine whether an object has been successfully removed, it fails to capture cases where negation modifies an entity’s state (e.g., a dog not running) or its properties (e.g., a not blue sphere). To address this limitation, we introduce more diverse prompts that encompass negation in objects, actions, and attributes as shown in Tab. 11).
Since object detectors are insufficient for evaluating these more complex negation cases, we drew inspiration from the TIFA [15] metric’s MLLM-based VQA system and introduce the Neg Score as a more comprehensive evaluation measure in Sec 5.2.
Appendix G Additional Qualitative Results